Approximately 14 million Americans are addicted to drugs and alcohol. The federal government estimates that these addicts cost the U.S. economy $300 billion in medical expenses and lost productivity. Despite the enormous potential market, many biotech companies have shied away from funding research and development (R&D) initiatives to find a cure for drug and alcohol addiction. Your firm – Drug Abuse Sciences (DAS) – is a notable exception. It has spent $280 million to date working on a cure, but is now at a crossroads. It can either abandon its program or invest another $75 million today. Unfortunately, the firm’s opportunity cost of funds is 8 percent and it will take another five years before final approval from the Federal Drug Administration is achieved and the product is actually sold. Expected (year-end) profits from selling the drug are presented in the accompanying table.

Year 1

Year 2

Year 3

Year 4

Year 5

Year 6

Year 7

Year 8

Year 9

$0

$0

$0

$0

$15,400,000

$17,500,000

$19,100,000

$22,700,000

$25,900,000

What is the net present value of the project?

Instruction: Enter your response rounded to the nearest penny (two decimal places). Use a negative sign (-) where appropriate.

$

Should DAS continue with its plan to bring the drug to market, or should it abandon the project?
Answer #1

P = F / (1+i)^t

Value invested until today is sunk cost and is not used in economic analysis

NPW of the project = -75000000 + 15400000 / (1+0.08)^5 + 17500000 / (1+0.08)^6 + 19100000 / (1+0.08)^7 + 22700000 / (1+0.08)^8 + 25900000 / (1+0.08)^9

= -75000000 + 15400000 / (1.08)^5 + 17500000 / (1.08)^6 + 19100000 / (1.08)^7 + 22700000 / (1.08)^8 + 25900000 / (1.08)^9

= -17125831.82

As NPV of project is negative, DAS should not continue this project